OpenStudy (anonymous):

1/4,15%,0.08,9.3% order from least to greatest

OpenStudy (welshfella):

i suggest you convert to percentages then it will be easy to compare 1/4 * 100 = ? 0.08 * 100 = ? can you calculate these?

OpenStudy (anonymous):

.08 = .080 9.3% = .093 15% = .150 1/4 = .250

OpenStudy (anonymous):

multiply your non-percentages 1/4 and 0.08 by 100 the answers you have are not correct

OpenStudy (welshfella):

yes 15% and 9.3% are percentages already

OpenStudy (anonymous):

look im in a rush and i need the answer not being rude just the truth

OpenStudy (welshfella):

that is why i only gave you only 2 calculations

OpenStudy (welshfella):

0.08 *100 - move the decimal poin 2 places to the right 1/4 = 0.25 - do same thing with decimal point

OpenStudy (anonymous):

oh ok

OpenStudy (welshfella):

when you find the correct order list them as in the question

OpenStudy (anonymous):

but i still dont get it

OpenStudy (anonymous):

@imqwerty may you help

imqwerty (imqwerty):

yeah sure :)

imqwerty (imqwerty):

1/4,15%,0.08,9.3% so we have all these^ \(\large \frac{1}{4}=0.25\) 15%\( =15 \times \large \frac{1}{100} =\frac{15}{100}=0.15\) and \(0.08\) is \(0.08\) 9.3% \( = 9.3 \times \large \frac{1}{100} = \frac{9.3}{100} =0.093\) so now you have->\(0.25,~0.15,~0.08,~0.093\) :) and now you can easily see which one is is big and which is small

imqwerty (imqwerty):

did you get the answer? i can help if you are stuck somewhere :)

OpenStudy (anonymous):

yes i did i medal you

imqwerty (imqwerty):

okay :)